# Self-Ensemble: Mitigating Confidence Distortion for Large Language Models

### v1

@@ 6 Related Work
Calibration aligns predicted probabilities with true correctness likelihoods to ensure reliable confidence estimates. Traditional methods like vector scaling Guo et al. (2017), isotonic regression Zadrozny and Elkan (2002), and Dirichlet calibration Kull et al. (2019) have been adapted to LLMs but often assume fixed prediction heads and struggle with distribution shifts or multi-choice uncertainty. Recent generative-specific techniques, such as logit perturbation Jiang et al. (2021), prompt-based reweighting Zhou et al. (2022b), and Bayesian post-hoc methods Wang et al. (2023), partially address these limitations. However, consistent calibration across diverse LLM architectures and tasks remains challenging, highlighting the need for effective, model-agnostic solutions.
@@
Ensembling has long been a powerful strategy to improve both model accuracy and robustness. In LLMs, methods such as voting schemes Zhou et al. (2022a), stochastic decoding Ho and Weller (2022), and diverse prompting Wang et al. (2022) aggregate outputs from multiple model instances or inference runs. These methods are especially useful for reducing variance in generative outputs and mitigating individual model biases. Additionally, ensemble approaches support calibration via confidence averaging and self-consistency mechanisms Wang et al. (2022). However, standard ensembles are computationally costly and less generalizable across model families and tasks, motivating for a more efficient and scalable ensemble method.
@@ 7 Conclusion
In this work, we demonstrate the confidence distortion problem of LLMs generally on MCQA, particularly in the many-choice setting. To solve this problem, we propose Self-Ensemble by decomposing a many-choice problem into several few-choice problems, and aggregating the intermediate results into the final solution. We further integrate Self-Ensemble with existing LLM architecture, enabling an intrinsic self-ensemble process for LLM inference. Experimental results across various LLMs and datasets show that Self-Ensemble can effectively overcome this problem, improving confidence in correct answers and reducing confidence in incorrect answers. This enables Self-Ensemble to significantly improve the accuracy of LLMs in MCQA by $8\%$ on average.
@@ 8 Limitations and Potential Risks
Appendix A General-purpose LLMs with Self-Ensemble Outperforms Reasoning LLMs
We show that general-purpose LLMs, when combined with Self-Ensemble, can surpass specialized reasoning LLMs on MCQA. Deepseek-R1-Distill models Guo et al. (2025) are distilled LLMs that internalize reasoning by fine-tuning on retrieved traces from Deepseek-R1. As shown in Table 5, both LLaMA-3-8B and Qwen2-7B, after applying Self-Ensemble, achieve higher accuracy compared to their reasoning-tuned counterparts Deepseek-R1-Distill-Llama-8B and Deepseek-R1-Distill-Qwen-7B on the QASC dataset. This highlights that Self-Ensemble can unlock the potential of general LLMs on MCQA without requiring dedicated reasoning fine-tuning.
Appendix B Self-Ensemble Outperforms Chain of Thought Prompting
We demonstrate that Self-Ensemble yields higher improvement on the QASC dataset than chain of thought (CoT) prompting Wei et al. (2022). CoT elicits intermediate reasoning steps from an LLM by prepending a prompt such as “Let’s think step by step” before generating the final answer to improve accuracy on MCQA. As shown in Table 5, with Self-Ensemble, both LLaMA-3-8B and Qwen2-7B models achieve higher accuracy on the QASC dataset than CoT prompting. While CoT prompting incurs additional inference latency and token-generation overhead by producing reasoning chains, Self-Ensemble shows its potential as a more efficient and effective calibration method on MCQA, requiring only a single forward pass with marginal overhead.
Appendix C Packages
In this work, we use the transformers along with datasets packages Wolf et al. (2019) for model and dataset loading. All open-sourced packages have the Apache-2.0 license, which allows for academic research.
@@ Appendix C Packages
LLaMA-3-8B 73.22
DeepSeek-R1-Distill-Llama-8B 73.33
LLaMA-3-8B with CoT 74.62
LLaMA-3-8B with Self-Ensemble 80.67
Qwen2-7B 69.44
DeepSeek-R1-Distill-Qwen-7B 70.63
Qwen2-7B with CoT 74.19
Qwen2-7B with Self-Ensemble 77.86
@@
Appendix D Computational Infrastructure

### v2

@@ 6 Related Work
Calibration aligns predicted probabilities with true correctness likelihoods to ensure reliable confidence estimates. Traditional methods like vector scaling Guo et al. (2017), isotonic regression Zadrozny and Elkan (2002), and Dirichlet calibration Kull et al. (2019) have been adapted to LLMs but often assume fixed prediction heads and struggle with distribution shifts or multi-choice uncertainty. Recent generative-specific techniques, such as logit perturbation Jiang et al. (2021), prompt-based reweighting Zhou et al. (2022b), Bayesian post-hoc methods Wang et al. (2023), chain-of-though reasoning Seßler et al. (2024) partially address these limitations. Moreover, pioneer work Zhong et al. (2025) proposes a unified framework of calibrating LLMs under weight quantization, with theoretical foundations and practical design. However, consistent calibration across diverse LLM architectures and tasks remains challenging, highlighting the need for effective, model-agnostic solutions.
@@
Ensembling has long been a powerful strategy to improve both model accuracy and robustness. In LLMs, methods such as voting schemes Zhou et al. (2022a), stochastic decoding Ho and Weller (2022), and diverse prompting Wang et al. (2022) aggregate outputs from multiple model instances or inference runs. Notably, similar to our intuition of self-ensemble, a concurrent work Neeley et al. (2025) also proposes a "divide-and-conquer" strategy to improve LLM performance on healthcare tasks. It has shown promising results in diagnosing rare diseases caused by genetic variants, a task that remains highly challenging. These methods are especially useful for reducing variance in generative outputs and mitigating individual model biases. Additionally, ensemble approaches support calibration via confidence averaging and self-consistency mechanisms Wang et al. (2022). However, standard ensembles are computationally costly and less generalizable across model families and tasks, motivating for a more efficient and scalable ensemble method.
@@ 7 Conclusion
In this work, we demonstrate the confidence mis-calibration of LLMs generally on MCQA, particularly in the many-choice setting. To solve this problem, we propose Self-Ensemble by decomposing a many-choice problem into several few-choice problems, and aggregating the intermediate results into the final solution. We further integrate Self-Ensemble with existing LLM architecture, enabling an intrinsic self-ensemble process for LLM inference. Experimental results across various LLMs and datasets show that Self-Ensemble can effectively overcome this problem, improving confidence in correct answers and reducing confidence in incorrect answers. This enables Self-Ensemble to significantly improve the accuracy of LLMs in MCQA by $8\%$ on average.
@@ 8 Limitations and Potential Risks
Acknowledgments
This research was supported by NSF CNS 2528780. We would like to thank The Center for Research Computing at Rice University for providing hardware and technical assistance for experiments. We extend special thanks to Xinyu Wang from the University of Kansas for help with the paper’s visualizations. The views and conclusions in this paper are those of the authors and do not represent the views of any funding or supporting agencies.
@@ Acknowledgments
Appendix A Efficiency Comparison of Multi-Pass and Single-Pass Self-Ensemble
We follow existing work Wang et al. (2024b); Liu et al. (2023); Yuan et al. (2024) to benchmark the efficiency of multi-pass Self-Ensemble in Section 3.2 and the single-pass Self-Ensemble in Section 4 using LLaMA-3-8B on the QASC dataset. As shown in Table 7, single-pass Self-Ensemble successfully accelerates inference time without sacrificing accuracy while maintaining almost the same memory cost. This demonstrates the effectiveness of our Section 4 in accelerating the inference through parallelization.
Appendix B Confidence Mis-calibration Across Model Scales and Families
We assess generality beyond 7B–8B models by evaluating LLaMA-3.2-1B/3B, LLaMA-3-8B, and Qwen2.5-1.5B/3B/7B/14B on QASC with $K\in\{2,4,6,8\}$. As shown in Table 8, accuracy consistently declines as $K$ increases for every model, indicating that the confidence mis-calibration phenomenon holds across architectures and scales.
Appendix C Extended Confidence Mis-calibration Evaluation
To further illustrate Self-Ensemble mitigates confidence mis-calibration, we provide additional evaluations based on the under-confidence and over-confidence ratios. Here, under-confidence is defined as the proportion of correct answers where the model assigns probability below a threshold $\tau$, while over-confidence is the proportion of incorrect answers where the model assigns probability above $\tau$. For example, with $\tau=0.5$, the under-confidence ratio measures the percentage of correct answers where $f(Q)<0.5$, and the over-confidence ratio measures the percentage of incorrect answers where $f(Q)>0.5$.
We report results averaged across LLaMA-3-8B, Mistral-7B-v0.1, and Qwen2-7B on QASC, TruthfulQA, and MMLU-Pro. As shown in Table 9, Self-Ensemble consistently reduces both under-confidence and over-confidence compared to standard inference, demonstrating its ability to mitigate confidence mis-calibration rather than exacerbate it.
Multi-pass Single-pass
Time (s) per question $\downarrow$ 0.78 0.16
Memory (GB) 14.97 15.12
Accuracy (%) 78.62 80.67
Table 7: Efficiency and accuracy comparison of multi-pass vs. single-pass Self-Ensemble
Model 2-choice 4-choice 6-choice 8-choice
LLaMA-3.2-1B 81.75 66.85 51.51 50.22
LLaMA-3.2-3B 91.04 84.77 77.97 69.55
LLaMA-3-8B 91.90 86.29 76.35 73.22
Qwen2.5-1.5B 89.30 81.53 62.31 60.91
Qwen2.5-3B 91.47 81.53 75.81 69.98
Qwen2.5-7B 94.71 89.42 82.51 76.57
Qwen2.5-14B 95.68 91.68 86.29 79.59
Appendix D Additional Self-Ensemble Results Across Model Scales
To assess the generality of Self-Ensemble, we report 8-choice QASC accuracy across two model families spanning multiple scales. As shown in Table 10, Self-Ensemble consistently improves over standard inference from small (1–3B) to larger (7–14B) models.
Under-confidence $\downarrow$
Model Standard Self-Ensemble
LLaMA-3-8B 43.45 35.33
Mistral-7B-v0.1 58.33 50.60
Qwen2-7B 41.41 34.25
Over-confidence $\downarrow$
Model Standard Self-Ensemble
LLaMA-3-8B 25.09 20.53
Mistral-7B-v0.1 33.77 21.66
Qwen2-7B 30.89 25.13
Table 9: Under-confidence and over-confidence ratios at $\tau=0.5$.
Appendix E Robustness to Random Grouping Seeds
We quantify the effect of grouping randomness by repeating Self-Ensemble on QASC with seeds $s\in\{0,1,2,3,4\}$ that control the random partitioning of answer choices, holding all other variables fixed. Accuracy is reported as mean $\pm$ standard deviation across seeds. As shown in Table 11, the small variances (all $\leq 0.60$ percentage points) indicate that Self-Ensemble is stable to grouping randomness across models.
Model Standard Self-Ensemble
Qwen2.5 family
Qwen2.5-1.5B 60.91 71.81
Qwen2.5-3B 69.98 76.78
Qwen2.5-7B 76.57 81.43
Qwen2.5-14B 79.59 83.37
LLaMA-3.x family
LLaMA-3.2-1B 50.22 56.37
LLaMA-3.2-3B 69.55 77.32
LLaMA-3-8B 73.22 80.67
Table 10: Self-Ensemble consistently improves accuracy over standard inference across scales.
Appendix F Group-wise Normalization with a Null Option
Self-Ensemble normalizes probabilities within each group by augmenting the group with a null choice ("None of the above") and applying a softmax over the augmented set. This guarantees per-group normalization (the probabilities over all candidates in a group sum to 1). Two immediate consequences explain the observed behavior: (i) when the correct answer is present, the null option receives negligible mass, so the sum over real options is close to 1; (ii) when the correct answer is absent, the null option absorbs probability mass, so the sum over real options drops below 1.
@@ Appendix F Group-wise Normalization with a Null Option
LLaMA-3-8B 80.24 $\pm$ 0.60
Mistral-7B-v0.1 77.91 $\pm$ 0.60
Qwen2-7B 65.87 $\pm$ 0.36
Table 11: Accuracy with error bar of Self-Ensemble on QASC across random seeds.
Appendix G Combination with Improvement
Self-Ensemble is designed to be orthogonal to existing approaches and can be seamlessly combined with them to further enhance the performance and reliability. Specifically, it can be integrated with efficient inference methods, such as KIVI Liu et al. (2024); Yuan et al. (2024) or H2O Zhang et al. (2023), to accelerate model execution while maintaining accuracy. Moreover, it can also be combined with secure deployment techniques, such as Taylor-Unswift Wang et al. (2024b), which effectively protects the intellectual property of LLMs through parameter space decomposition. Finally, it is compatible with methods for multi-agent frameworks for solving complex tasks, such as  Chang et al. (2024); Yu et al. (2025).
Appendix H Packages
In this work, we use the transformers along with datasets packages Wolf et al. (2019) for model and dataset loading. All open-sourced packages have the Apache-2.0 license, which allows for academic research. We use public benchmarks and open models strictly for research and evaluation in accordance with their original licenses/terms of use. We release Self-Ensemble code, prompts, and evaluation scripts for research reproducibility; any derivatives of licensed datasets inherit their original research-only restrictions.
@@ Appendix H Packages
Appendix I Computational Infrastructure